# Effective grand-canonical description of condensation in negative-temperature regimes

### v1

@@ Conclusions
Recent large-deviation studies, based on a microcanonical analysis of finite systems at infinite times, pointed out the existence of a first-order condensation transition at $\beta=0$, accompanied by relevant finite-$N$ corrections on the NT side. Here, we have revisited the problem by implementing a grand-canonical approach, thanks to the introduction of thermal reservoirs, which operate at large NTs. Our key result is the evidence of a temperature-dependent critical amplitude, separating stable from unstable condensation peaks (DNLS breathers). As a result, close to the $\beta=0$ line, extended states are metastable but characterized by an exponentially long lifetime. In the DNLS model, the exponential growth is enhanced by the fast decrease of the diffusion coefficient with the breather height. In this Letter we have assumed the presence of a single breather; in general, multiple breathers are expected to arise in the NT region, each accompanied by an additional quasi-conserved quantity (the AI). It is conceivable that a generalized Gibbs ensembles should be implemented, analogous to what done in the context of integrable systems Vidmar and Rigol (2016); Spohn (2020). This generalization is, however, not straightforward; it requires further work.
@@ Acknowledgements.
We thank P. Politi for fruitful discussions.

### v2

@@ Conclusions
We have revisited the stability of extended states by implementing a grand-canonical approach and introducing suitable thermal reservoirs, which operate at large NTs. Our main achievement is the demonstration of metastable states characterized by exponentially long lifetimes. This result follows from the existence of a temperature-dependent critical amplitude, which separates stable from unstable condensation peaks (DNLS breathers). In the DNLS model, the exponential growth is amplified by the fast decrease of the diffusion coefficient with the breather height.
In this Letter, we focused on a setup where the local onsite energy $\epsilon$ is a quadratic function of the mass $c$, but the formalism can be applied to contexts where $\epsilon(c)>0$ (see, e.g., purely stochastic condensation models Szavits-Nossan et al. (2014a, b)): in such cases $c_{max}$ solves $\beta\epsilon^{\prime}(c_{max})=m$. Finally, we have assumed the presence of a single breather; in general, multiple breathers are expected to arise in the NT region, each accompanied by an additional quasi-conserved quantity. It is conceivable that a generalized Gibbs ensemble should be implemented, analogous to what done in the context of integrable systems Vidmar and Rigol (2016); Spohn (2020). This generalization is, however, not straightforward; it requires further work.
@@ Acknowledgements.
We thank P. Politi for fruitful discussions. We acknowledge support from the project “Breakdown of ergodicity in classical and quantum many-body systems” (BECQuMB) funded by the Italian Ministry for University and Research (MUR) PRIN2022 grant No. 20222BHC9Z.